CSePub Home Book StoreInstructors Independent AuthorsAbos Dagpurd Incorporated has a bond tradrg onte secondary maket hal wil mulrenfour yeal The bond paysananual coupon weacopon rle of6S0%ad has a face value of $1,00000 Based onthe economy and nsk assoated ㆋh Dagpund you seek a 1225% return on Dag ord debt vhat price are you wing to pay for the bord? Answer Format: Currency Round to: 2 decimal places Enter Answer Here Subnt Answer t10 15 i6 17
Answer #1

PVIFA for n=4 and i=12.5% = [1-(1+0.1225)^-4]/0.1225 = 3.02143

PV of Coupon payments = 65*3.02143 = $196.39

Annuity factor of n=4 and i=12.5% = 0.62988

PV of Maturity value = $1,000 * 0.62988 = $629.88

So, the bond price we will be willing to pay = $629.88 + $196.39 = $826.27